OpenStudy (anonymous):

Solve 9c + 4 = −23 −3 3 −2.1 4

OpenStudy (mathstudent55):

You need to isolate c. FIrst, subtract 4 from both sides. Then divide both sides by 9.

OpenStudy (anonymous):

9c + 4 = −23 SUBTRACT 4 FROM BOTH SIDES 9c + 4 = −23 -4 -4 ---- ---- WHAT DO YOU GET?

OpenStudy (anonymous):

-19

OpenStudy (mathstudent55):

-23 - 4 = -27

OpenStudy (anonymous):

ok so now u have 9c=-27 no u have to divide both sides by 9 -27 divided by 9 c =-3

OpenStudy (anonymous):

^ let the user answer. do not just give out the answer without letting the user answer first. thank you

OpenStudy (anonymous):

i am not just giving the answer i am also helping